The relativistic kinetic energy of a particle is three times its rest mass energy. Find
the:
a) Lorentz factor and speed parameter
b) velocity (in m/s) of the particle.
1
Expert's answer
2020-05-18T10:22:07-0400

Given data

Relativistic Kinetic energy of particle is "K=3E_0"

Here "E_0"  is the rest energy of particle.

a)

The expression for relativistic kinetic energy of particle is given by

"K=(\gamma-1)E_0"

Here "\gamma =\frac{1}{\sqrt{\beta^2-1}}" is called relativistic factor or Lorentz factor.

"3E_0=(\\gamma-1)E_0 \u200b"

"\gamma=4"

Hence the Lorentz factor is "\gamma=4" .

But the Lorentz factor is

"\gamma =\frac{1}{\sqrt{\beta^2-1}}"

"4=\\frac{1}{\\sqrt{\\beta^2-1}} \u200b"

Solve for "\beta" ,

"\beta =0.968"

So the velocity factor is"\u03b2=0.968."

----------------------------------------------------------------------------------

b)

The velocity factor is

"\u03b2=0.968"

"\frac{v}{c}=0.968"

Therefore, the velocity of the particle is

"v=0.968c=0.968(3*10^8 m\/s)=2.90*10^8 m\/s"